### v1

@@ 1 Introduction
This paper aims to answer a long-time open problem in variational analysis: whether the Aubin property of the solution mapping associated with the canonically perturbed Karush-Kuhn-Tucker (KKT) system and the strong regularity of the KKT system are equivalent for nonlinear conic programming without convexity at a locally optimal solution. In the context of nonlinear SOCP [1], here we completely settle this question by proving the equivalence of the two concepts in the absence of strict complementarity. To be precise, we consider the nonlinear SOCP problem of the form
@@
Note that $\mbox{({\bf ii})}\Rightarrow\mbox{({\bf i})}$ trivially holds. Moreover, according to [26, Corollary 2.2], $({\bf ii})$ is equivalent to the condition that $({\boldsymbol{x}}^{*},{\boldsymbol{y}}^{*})$ is a strongly regular solution (in the sense of Robinson [26]) of the generalized equation
@@
i.e., there exist two neighborhoods ${\mathcal{U}}$ of the origin $({\boldsymbol{0}},{\boldsymbol{0}})\in{\mathcal{X}}\times{\mathcal{Y}}$ and ${\mathcal{V}}$ of $({\boldsymbol{x}}^{*},{\boldsymbol{y}}^{*})$, respectively, such that for every $({\boldsymbol{a}},{\boldsymbol{b}})\in{\mathcal{U}}$, the linearized generalized equation (c.f. [5, Definition 22] or [23, Definition 13]) has a unique solution in ${\mathcal{V}}$, denoted by $q_{{\mathcal{V}}}({\boldsymbol{a}},{\boldsymbol{b}})$, and the mapping $q_{{\mathcal{V}}}:{\mathcal{U}}\to{\mathcal{V}}$ is locally Lipschitz continuous. Since the generalized equation (1.6) is a reformulation of the KKT system (1.4) without perturbation, a strongly regular solution to (1.6) is also called a strongly regular solution to the KKT system of (1.2).
@@
where ${\sf lin}{\mathcal{T}}_{{\mathcal{K}}}(G({\boldsymbol{x}}))$ denotes the linearity space, i.e. the largest linear space contained in this tangent cone (in the sense of convex analysis). The original definition of constrain nondegeneracy takes the form of [3, Definition 2.1], and (1.7) is an equivalent reformulation given in [5, Definition 16] since ${\mathcal{K}}$ is $C^{2}$-reducible to a closed convex cone [5, Lemma 15]. In particular, it is well-known [5, Theorem 30] that $({\bf ii})$ is equivalent to the strong second-order sufficient condition of SOCP (1.1) at ${\boldsymbol{x}}^{*}$ (c.f. (4.6)) and the constraint nondegeneracy. Moreover, these two conditions have played a key role in proving the fast linear or superlinear local convergence rates of augmented Lagrangian methods [7, 16, 11, 30]. Additionally, the strong regularity condition also implies the full stability of locally optimal solutions for (1.2). One may refer to [19, 21] for more information.
@@
Under the assumption that the strict complementarity condition holds except for possibly only one block of the constraints, the equivalence of (i) and (ii) was proved in [23] and [22]. As was concluded in [23, Section 5], the authors were uncertain whether such a restriction should be attributed to their proof technique based on the S-lemma, or the intrinsic property of the second-order cones. Note that, however, for convex programming, the equivalence between the two concepts is known from [9, Proposition 5.1].
@@ 2 Notation and preliminaries
It is easy to see that Assertion (2) is not true if Assertion (1) holds. On the other hand, suppose that the Assertion (1) does not hold, i.e., $\exists\,{\boldsymbol{u}}\in{\sf int}Q$ such that $\langle{\boldsymbol{u}},{\mathcal{M}}{\boldsymbol{u}}\rangle\leq 0$. Consider the following optimization problem
@@
Since the feasible set of (2.1) is compact and $\langle{\boldsymbol{u}},{\mathcal{M}}{\boldsymbol{u}}\rangle>0$, $\forall\,{\boldsymbol{u}}\in{\sf bdry}Q\setminus\{{\boldsymbol{0}}\}$, there exists a ${\boldsymbol{u}}^{*}\in{\sf int}Q$ globally solves (2.1) with $\langle{\boldsymbol{u}}^{*},{\mathcal{M}}{\boldsymbol{u}}^{*}\rangle\leq 0$. It is also easy to check that Robinson’s constraint qualification holds at ${\boldsymbol{u}}^{*}$ for (2.1). Consequently, there exists a multiplier $({\boldsymbol{y}},\lambda)\in{\mathcal{E}}\times{\mathds{R}}$, together with ${\boldsymbol{u}}^{*}$, satisfies
@@
The second lemma is a characterization of positive definite matrices originates from [13, Theorem 3.6], and one may refer to [14, Proposition 3.1] for a more straightforward proof based on Moreau decomposition.
@@
In the following, we prepare some definitions and preliminary results in variational analysis. The regular (Fréchet) normal cone to $\Omega$ at ${\boldsymbol{u}}\in\Omega$ is defined by
@@
According to [18, Theorem 1.6], the limiting (Mordukhovich) normal cone to $\Omega$ at ${\boldsymbol{u}}$ can be defined by
@@
Also, one has for any $v\in{\mathds{R}}$, it holds that
@@
According to [23, Lemma 19], we have the following result on coderivatives.
@@
Since ${\mathcal{K}}$ is ${\mathcal{C}}^{2}$-cone reducible at every ${\boldsymbol{y}}\in{\mathcal{K}}$, one can use the second-order chain rule developed in [23, Theorem 7], which is a generalization of [20, Theorem 3.4], to get the following result based on repeating the proof of [23, Theorem 20].
@@
Let ${\boldsymbol{x}}^{*}$ be a stationary point to (1.2). Suppose that the constraint nondegeneracy (1.7) holds at ${\boldsymbol{x}}^{*}$, and ${\boldsymbol{y}}^{*}$ is the corresponding multiplier. Then one has
@@ 3 Coderivatives of the second-order cone
Let $r\geq 1$ and ${\mathcal{Q}}:=\{{\boldsymbol{z}}=(\dot{z};\bar{\boldsymbol{z}})\in{\mathds{R}}^{1+r}\mid\dot{z}\geq\|\bar{\boldsymbol{z}}\|\}$ be a second-order cone. From [1, Lemma 15] we know that for any ${\boldsymbol{z}},{\boldsymbol{v}}\in{\mathcal{Q}}$ one has $\langle{\boldsymbol{z}},{\boldsymbol{v}}\rangle=0$ if and only if either ${\boldsymbol{z}}={\boldsymbol{0}}$ or ${\boldsymbol{v}}={\boldsymbol{0}}$, or there exists a scalar $\alpha>0$ such that ${\boldsymbol{z}}=\alpha(\dot{v};-\bar{\boldsymbol{v}})$. For a given vector ${\boldsymbol{z}}\in{\mathds{R}}^{1+r}$, its spectral factorization associated with ${\mathcal{Q}}$ is given by ${\boldsymbol{z}}=\sigma_{1}({\boldsymbol{z}}){\boldsymbol{c}}^{1}({\boldsymbol{z}})+\sigma_{2}({\boldsymbol{z}}){\boldsymbol{c}}^{2}({\boldsymbol{z}})$, where
@@
with ${\boldsymbol{w}}\in\mathbb{R}^{r}$ being an arbitrarily vector such that $\|{\boldsymbol{w}}\|=1$. Then, according to [12, Proposition 3.3], the explicit formula of $\Pi_{{\mathcal{Q}}}$ can be given as
@@
Moreover, the explicit formula of $\partial_{B}\Pi_{{\mathcal{Q}}}$ has been computed in [25, Lemma 14], [6, Lemma 4], and [15, Lemmas 2.5 & 2.6]. In particular, one has
@@
$\begin{array}[]{ll}W{\boldsymbol{p}}=(I-W){\boldsymbol{q}}\\[2.84526pt] \Leftrightarrow\dot{p}+\frac{\bar{\boldsymbol{u}}^{\top}\bar{\boldsymbol{p}}}{\|\bar{\boldsymbol{u}}\|}=0\ \mbox{and}\ -\frac{\bar{\boldsymbol{u}}^{\top}\bar{\boldsymbol{p}}}{\|\bar{\boldsymbol{u}}\|^{2}}\bar{\boldsymbol{u}}+\frac{2\dot{u}}{{\dot{u}}-{\dot{v}}}\bar{\boldsymbol{p}}-\frac{{\dot{u}}+{\dot{v}}}{{\dot{u}}-{\dot{v}}}\frac{\bar{\boldsymbol{u}}^{\top}\bar{\boldsymbol{p}}}{\|\bar{\boldsymbol{u}}\|^{2}}\bar{\boldsymbol{u}}=-\frac{\bar{\boldsymbol{u}}^{\top}\bar{\boldsymbol{q}}}{\|\bar{\boldsymbol{u}}\|^{2}}\bar{\boldsymbol{u}}-\frac{2\dot{v}}{{\dot{u}}-{\dot{v}}}\bar{\boldsymbol{q}}+\frac{{\dot{u}}+{\dot{v}}}{{\dot{u}}-{\dot{v}}}\frac{\bar{\boldsymbol{u}}^{\top}\bar{\boldsymbol{q}}}{\|\bar{\boldsymbol{u}}\|^{2}}\bar{\boldsymbol{u}}\\ \Leftrightarrow\dot{p}+\frac{\bar{\boldsymbol{u}}^{\top}\bar{\boldsymbol{p}}}{\|\bar{\boldsymbol{u}}\|}=0\ \mbox{and}\ -\frac{\bar{\boldsymbol{u}}\bar{\boldsymbol{u}}^{\top}}{\|\bar{\boldsymbol{u}}\|^{2}}\bar{\boldsymbol{p}}+\frac{2\dot{u}}{{\dot{u}}-{\dot{v}}}\bar{\boldsymbol{p}}-\frac{{\dot{u}}+{\dot{v}}}{{\dot{u}}-{\dot{v}}}\frac{\bar{\boldsymbol{u}}\bar{\boldsymbol{u}}^{\top}}{\|\bar{\boldsymbol{u}}\|^{2}}\bar{\boldsymbol{p}}=-\frac{\bar{\boldsymbol{u}}\bar{\boldsymbol{u}}^{\top}}{\|\bar{\boldsymbol{u}}\|^{2}}\bar{\boldsymbol{q}}-\frac{2\dot{v}}{{\dot{u}}-{\dot{v}}}\bar{\boldsymbol{q}}+\frac{{\dot{u}}+{\dot{v}}}{{\dot{u}}-{\dot{v}}}\frac{\bar{\boldsymbol{u}}\bar{\boldsymbol{u}}^{\top}}{\|\bar{\boldsymbol{u}}\|^{2}}\bar{\boldsymbol{q}}\\ \Leftrightarrow\dot{p}+\frac{\bar{\boldsymbol{u}}^{\top}\bar{\boldsymbol{p}}}{\|\bar{\boldsymbol{u}}\|}=0\ \mbox{and}\ \frac{2\dot{u}}{{\dot{u}}-{\dot{v}}}\bar{\boldsymbol{p}}-\frac{2{\dot{u}}}{{\dot{u}}-{\dot{v}}}\frac{\bar{\boldsymbol{u}}\bar{\boldsymbol{u}}^{\top}}{\|\bar{\boldsymbol{u}}\|^{2}}\bar{\boldsymbol{p}}=-\frac{2\dot{v}}{{\dot{u}}-{\dot{v}}}\bar{\boldsymbol{q}}+\frac{2{\dot{v}}}{{\dot{u}}-{\dot{v}}}\frac{\bar{\boldsymbol{u}}\bar{\boldsymbol{u}}^{\top}}{\|\bar{\boldsymbol{u}}\|^{2}}\bar{\boldsymbol{q}}\\ \Leftrightarrow\dot{p}+\frac{\bar{\boldsymbol{u}}^{\top}\bar{\boldsymbol{p}}}{\|\bar{\boldsymbol{u}}\|}=0\ \mbox{and}\ (I_{r}-\frac{\bar{\boldsymbol{u}}\bar{\boldsymbol{u}}^{\top}}{\|\bar{\boldsymbol{u}}\|^{2}})\bar{\boldsymbol{p}}=-\frac{\dot{v}}{\dot{u}}(I_{r}-\frac{\bar{\boldsymbol{u}}\bar{\boldsymbol{u}}^{\top}}{\|\bar{\boldsymbol{u}}\|^{2}})\bar{\boldsymbol{q}}\\ \Leftrightarrow\dot{p}+\frac{\bar{\boldsymbol{u}}^{\top}\bar{\boldsymbol{p}}}{\|\bar{\boldsymbol{u}}\|}=0\ \mbox{and}\ (I_{r}-\frac{\bar{\boldsymbol{u}}\bar{\boldsymbol{u}}^{\top}}{\|\bar{\boldsymbol{u}}\|^{2}})(\bar{\boldsymbol{p}}+\frac{\dot{v}}{\dot{u}}\bar{\boldsymbol{q}})=0\\ \Leftrightarrow\dot{p}+\frac{\bar{\boldsymbol{u}}^{\top}\bar{\boldsymbol{p}}}{\|\bar{\boldsymbol{u}}\|}=0\quad\mbox{and}\quad\bar{\boldsymbol{p}}\in\big{\{}\tau\frac{\bar{\boldsymbol{u}}}{\|\bar{\boldsymbol{u}}\|}-\frac{\dot{v}}{\dot{u}}\bar{\boldsymbol{q}}\mid\tau\in{\mathds{R}}\big{\}}\\[2.84526pt] \Leftrightarrow{\boldsymbol{p}}\in\left\{\left(\frac{\dot{v}}{\dot{u}}\dot{q}-\tau;\,-\frac{\dot{v}}{\dot{u}}\bar{\boldsymbol{q}}+\tau\frac{\bar{\boldsymbol{u}}}{\|\bar{\boldsymbol{u}}\|}\right)\ \mid\ \tau\in{\mathds{R}}\right\},\end{array}$

### v2

@@ 1 Introduction
This paper aims to answer a long-time open problem in variational analysis: whether the Aubin property of the solution mapping associated with the canonically perturbed Karush-Kuhn-Tucker (KKT) system and the strong regularity of the KKT system are equivalent for nonlinear conic programming without assuming convexity at a locally optimal solution. In the context of nonlinear SOCP [1], here we completely settle this question by proving the equivalence of the two concepts in the absence of strict complementarity. To be precise, we consider the nonlinear SOCP problem of the form
@@
Note that $\mbox{({\bf ii})}\Rightarrow\mbox{({\bf i})}$ trivially holds. Moreover, according to [30, Corollary 2.2], $({\bf ii})$ is equivalent to the condition that $({\boldsymbol{x}}^{*},{\boldsymbol{y}}^{*})$ is a strongly regular solution (in the sense of Robinson [30]) of the generalized equation
@@
i.e., there exist two neighborhoods ${\mathcal{U}}$ of the origin $({\boldsymbol{0}},{\boldsymbol{0}})\in{\mathcal{X}}\times{\mathcal{Y}}$ and ${\mathcal{V}}$ of $({\boldsymbol{x}}^{*},{\boldsymbol{y}}^{*})$, respectively, such that for every $({\boldsymbol{a}},{\boldsymbol{b}})\in{\mathcal{U}}$, the linearized generalized equation (c.f. [6, Definition 22] or [27, Definition 13]) has a unique solution in ${\mathcal{V}}$, denoted by $q_{{\mathcal{V}}}({\boldsymbol{a}},{\boldsymbol{b}})$, and the mapping $q_{{\mathcal{V}}}:{\mathcal{U}}\to{\mathcal{V}}$ is locally Lipschitz continuous. Since the generalized equation (1.6) is a reformulation of the KKT system (1.4) without perturbation, a strongly regular solution to (1.6) is also called a strongly regular solution to the KKT system of (1.2).
@@
where ${\sf lin}{\mathcal{T}}_{{\mathcal{K}}}(G({\boldsymbol{x}}))$ denotes the linearity space, i.e. the largest linear space contained in this tangent cone (in the sense of convex analysis). The original definition of constrain nondegeneracy takes the form of [4, Definition 2.1], and (1.7) is an equivalent reformulation given in [6, Definition 16] since ${\mathcal{K}}$ is $C^{2}$-reducible to a closed convex cone [6, Lemma 15]. In particular, it is well-known [6, Theorem 30] that $({\bf ii})$ is equivalent to the strong second-order sufficient condition of SOCP (1.1) at ${\boldsymbol{x}}^{*}$ (c.f. (4.6)) and the constraint nondegeneracy. Moreover, these two conditions have played a key role in proving the fast linear or superlinear local convergence rates of augmented Lagrangian methods [9, 19, 13, 25, 36]. Additionally, the strong regularity condition also implies the full stability of locally optimal solutions for (1.2). One may refer to [22, 24] for more information.
@@
Under the assumption that the strict complementarity condition holds except for possibly only one block of the constraints, the equivalence of (i) and (ii) was proved in [27] and [26]. As was concluded in [27, Section 5], the authors were uncertain whether such a restriction should be attributed to their proof technique based on the S-lemma, or the intrinsic property of the second-order cones. Note that, however, for convex programming, the equivalence between the two concepts is known from [11, Proposition 5.1]. In a more general problem stetting, including problem (1.2) with ${\mathcal{K}}$ being an arbitrary closed convex set, a recent work [3, Theorem 4.2] shows that the variational sufficient condition for local optimality [31] at ${\boldsymbol{x}}^{*}$ is a sufficient condition for the equivalence between (i) and (ii). The restrictiveness of this variational sufficient condition will be discussed in Remark 4.1. Moreover, if additionally ${\mathcal{K}}$ is a $C^{2}$-cone reducible set, another recent work [16, Theorem 5.14] utilized an assumption on relative interiors of subdifferential mappings to prove the equivalence between (i) and (ii), and such a condition becomes the strict complementarity condition if (1.1) is the problem in consideration.
@@ 2 Notation and preliminaries
It is easy to see that Assertion (2) is not true if Assertion (1) holds. On the other hand, suppose that Assertion (1) does not hold, i.e., $\exists\,{\boldsymbol{u}}\in{\sf int}Q$ such that $\langle{\boldsymbol{u}},{\mathcal{M}}{\boldsymbol{u}}\rangle\leq 0$. Consider the following optimization problem
@@
Since the feasible set of (2.1) is compact and $\langle{\boldsymbol{u}},{\mathcal{M}}{\boldsymbol{u}}\rangle>0$, $\forall\,{\boldsymbol{u}}\in{\sf bdry}Q\setminus\{{\boldsymbol{0}}\}$, there exists a ${\boldsymbol{u}}^{*}\in{\sf int}Q$ that globally solves (2.1) with $\langle{\boldsymbol{u}}^{*},{\mathcal{M}}{\boldsymbol{u}}^{*}\rangle\leq 0$. It is also easy to check that Robinson’s constraint qualification [5, Equation (2.182)] holds at ${\boldsymbol{u}}^{*}$ for (2.1). Consequently, there exists a multiplier $({\boldsymbol{y}},\lambda)\in{\mathcal{E}}\times{\mathds{R}}$, together with ${\boldsymbol{u}}^{*}$, satisfying
@@
The second lemma is a characterization of positive definite matrices that originates from [15, Theorem 3.6], and one may refer to [17, Proposition 3.1] for a more straightforward proof based on the Moreau decomposition.
@@
In the following, we present some definitions and preliminary results in variational analysis. The regular (Fréchet) normal cone to $\Omega$ at ${\boldsymbol{u}}\in\Omega$ is defined by
@@
According to [21, Theorem 1.6], the limiting (Mordukhovich) normal cone to $\Omega$ at ${\boldsymbol{u}}$ can be defined by
@@
Also, for any $v\in{\mathds{R}}$ one has
@@
According to [27, Lemma 19], we have the following result on coderivatives.
@@
Since ${\mathcal{K}}$ is ${\mathcal{C}}^{2}$-cone reducible at every ${\boldsymbol{y}}\in{\mathcal{K}}$, one can use the second-order chain rule developed in [27, Theorem 7], which is a generalization of [23, Theorem 3.4], to get the following result based on repeating the proof of [27, Theorem 20].
@@
Let ${\boldsymbol{x}}^{*}$ be a stationary point to (1.2). Suppose that the constraint nondegeneracy (1.7) holds at ${\boldsymbol{x}}^{*}$, and ${\boldsymbol{y}}^{*}$ is the corresponding multiplier. Then for the mapping ${{\mathbb{S}}}_{\rm GE}$ defined by (1.8) one has
@@ 3 Coderivatives of the second-order cone
Let $r\geq 1$ be an integer and ${\mathcal{Q}}:=\{{\boldsymbol{z}}=(\dot{z};\bar{\boldsymbol{z}})\in{\mathds{R}}^{1+r}\mid\dot{z}\geq\|\bar{\boldsymbol{z}}\|\}$ be a second-order cone. From [1, Lemma 15] we know that for any ${\boldsymbol{z}},{\boldsymbol{v}}\in{\mathcal{Q}}$ one has $\langle{\boldsymbol{z}},{\boldsymbol{v}}\rangle=0$ if and only if either ${\boldsymbol{z}}={\boldsymbol{0}}$ or ${\boldsymbol{v}}={\boldsymbol{0}}$, or there exists a scalar $\alpha>0$ such that ${\boldsymbol{z}}=\alpha(\dot{v};-\bar{\boldsymbol{v}})$. For a given vector ${\boldsymbol{z}}\in{\mathds{R}}^{1+r}$, its spectral factorization associated with ${\mathcal{Q}}$ is given by ${\boldsymbol{z}}=\sigma_{1}({\boldsymbol{z}}){\boldsymbol{c}}^{1}({\boldsymbol{z}})+\sigma_{2}({\boldsymbol{z}}){\boldsymbol{c}}^{2}({\boldsymbol{z}})$, where
@@
with ${\boldsymbol{w}}\in\mathbb{R}^{r}$ being an arbitrarily vector such that $\|{\boldsymbol{w}}\|=1$. Then, according to [14, Proposition 3.3], the explicit formula of $\Pi_{{\mathcal{Q}}}$ can be given as
@@
Moreover, the explicit formula of $\partial_{B}\Pi_{{\mathcal{Q}}}$ has been computed in [29, Lemma 14], [7, Lemma 4], and [18, Lemmas 2.5 & 2.6]. In particular, one has
@@
$\begin{array}[]{ll}W{\boldsymbol{p}}=(I-W){\boldsymbol{q}}\\[2.84526pt] \Leftrightarrow\dot{p}+\frac{\bar{\boldsymbol{u}}^{\top}\bar{\boldsymbol{p}}}{\|\bar{\boldsymbol{u}}\|}=0\ \mbox{and}\ -\frac{\bar{\boldsymbol{u}}^{\top}\bar{\boldsymbol{p}}}{\|\bar{\boldsymbol{u}}\|^{2}}\bar{\boldsymbol{u}}+\frac{2\dot{u}}{{\dot{u}}-{\dot{v}}}\bar{\boldsymbol{p}}-\frac{{\dot{u}}+{\dot{v}}}{{\dot{u}}-{\dot{v}}}\frac{\bar{\boldsymbol{u}}^{\top}\bar{\boldsymbol{p}}}{\|\bar{\boldsymbol{u}}\|^{2}}\bar{\boldsymbol{u}}=-\frac{\bar{\boldsymbol{u}}^{\top}\bar{\boldsymbol{q}}}{\|\bar{\boldsymbol{u}}\|^{2}}\bar{\boldsymbol{u}}-\frac{2\dot{v}}{{\dot{u}}-{\dot{v}}}\bar{\boldsymbol{q}}+\frac{{\dot{u}}+{\dot{v}}}{{\dot{u}}-{\dot{v}}}\frac{\bar{\boldsymbol{u}}^{\top}\bar{\boldsymbol{q}}}{\|\bar{\boldsymbol{u}}\|^{2}}\bar{\boldsymbol{u}}\\ \Leftrightarrow\dot{p}+\frac{\bar{\boldsymbol{u}}^{\top}\bar{\boldsymbol{p}}}{\|\bar{\boldsymbol{u}}\|}=0\ \mbox{and}\ -\frac{\bar{\boldsymbol{u}}\bar{\boldsymbol{u}}^{\top}}{\|\bar{\boldsymbol{u}}\|^{2}}\bar{\boldsymbol{p}}+\frac{2\dot{u}}{{\dot{u}}-{\dot{v}}}\bar{\boldsymbol{p}}-\frac{{\dot{u}}+{\dot{v}}}{{\dot{u}}-{\dot{v}}}\frac{\bar{\boldsymbol{u}}\bar{\boldsymbol{u}}^{\top}}{\|\bar{\boldsymbol{u}}\|^{2}}\bar{\boldsymbol{p}}=-\frac{\bar{\boldsymbol{u}}\bar{\boldsymbol{u}}^{\top}}{\|\bar{\boldsymbol{u}}\|^{2}}\bar{\boldsymbol{q}}-\frac{2\dot{v}}{{\dot{u}}-{\dot{v}}}\bar{\boldsymbol{q}}+\frac{{\dot{u}}+{\dot{v}}}{{\dot{u}}-{\dot{v}}}\frac{\bar{\boldsymbol{u}}\bar{\boldsymbol{u}}^{\top}}{\|\bar{\boldsymbol{u}}\|^{2}}\bar{\boldsymbol{q}}\\ \Leftrightarrow\dot{p}+\frac{\bar{\boldsymbol{u}}^{\top}\bar{\boldsymbol{p}}}{\|\bar{\boldsymbol{u}}\|}=0\ \mbox{and}\ \frac{2\dot{u}}{{\dot{u}}-{\dot{v}}}\bar{\boldsymbol{p}}-\frac{2{\dot{u}}}{{\dot{u}}-{\dot{v}}}\frac{\bar{\boldsymbol{u}}\bar{\boldsymbol{u}}^{\top}}{\|\bar{\boldsymbol{u}}\|^{2}}\bar{\boldsymbol{p}}=-\frac{2\dot{v}}{{\dot{u}}-{\dot{v}}}\bar{\boldsymbol{q}}+\frac{2{\dot{v}}}{{\dot{u}}-{\dot{v}}}\frac{\bar{\boldsymbol{u}}\bar{\boldsymbol{u}}^{\top}}{\|\bar{\boldsymbol{u}}\|^{2}}\bar{\boldsymbol{q}}\\ \Leftrightarrow\dot{p}+\frac{\bar{\boldsymbol{u}}^{\top}\bar{\boldsymbol{p}}}{\|\bar{\boldsymbol{u}}\|}=0\ \mbox{and}\ (I_{r}-\frac{\bar{\boldsymbol{u}}\bar{\boldsymbol{u}}^{\top}}{\|\bar{\boldsymbol{u}}\|^{2}})\bar{\boldsymbol{p}}=-\frac{\dot{v}}{\dot{u}}(I_{r}-\frac{\bar{\boldsymbol{u}}\bar{\boldsymbol{u}}^{\top}}{\|\bar{\boldsymbol{u}}\|^{2}})\bar{\boldsymbol{q}}\\ \Leftrightarrow\dot{p}+\frac{\bar{\boldsymbol{u}}^{\top}\bar{\boldsymbol{p}}}{\|\bar{\boldsymbol{u}}\|}=0\ \mbox{and}\ (I_{r}-\frac{\bar{\boldsymbol{u}}\bar{\boldsymbol{u}}^{\top}}{\|\bar{\boldsymbol{u}}\|^{2}})(\bar{\boldsymbol{p}}+\frac{\dot{v}}{\dot{u}}\bar{\boldsymbol{q}})={{\boldsymbol{0}}}\\ \Leftrightarrow\dot{p}+\frac{\bar{\boldsymbol{u}}^{\top}\bar{\boldsymbol{p}}}{\|\bar{\boldsymbol{u}}\|}=0\quad\mbox{and}\quad\bar{\boldsymbol{p}}\in\big{\{}\tau\frac{\bar{\boldsymbol{u}}}{\|\bar{\boldsymbol{u}}\|}-\frac{\dot{v}}{\dot{u}}\bar{\boldsymbol{q}}\mid\tau\in{\mathds{R}}\big{\}}\\[2.84526pt] \Leftrightarrow{\boldsymbol{p}}\in\left\{\left(\frac{\dot{v}}{\dot{u}}\dot{q}-\tau;\,-\frac{\dot{v}}{\dot{u}}\bar{\boldsymbol{q}}+\tau\frac{\bar{\boldsymbol{u}}}{\|\bar{\boldsymbol{u}}\|}\right)\ \mid\ \tau\in{\mathds{R}}\right\},\end{array}$